A rectangular trench is to be dug which is to be 300 ft long, 36 ft wide and 10 ft deep. A large dump truck can carry 20 cubic yards in a load. How many truck loads will be required to remove the dirt from this trench?
Answer by stanbon(48568)

A rectangular trench is to be dug which is to be 300 ft long, 36 ft wide and 10 ft deep. A large dump truck can carry 20 cubic yards in a load. How many truck loads will be required to remove the dirt from this trench?
---------------------------------------
volume = (300*36*10) = 108000 cu. ft
each cubic yard = 3*3*3 = 27 cu ft
20 cu. yards = 20*27 cu ft. = 540 cu ft.
----------
# of truck loads = 108000/540 = 200

=============
Cheers,
Stan H.